OpenStudy (anonymous):

What is the inverse of g? g(x) = 4 - x

OpenStudy (abb0t):

to find the inverse, simply solve for "x" in terms of "y". So solve for x x = 4 - y

OpenStudy (jdoe0001):

@sshehan so, what did you get for \(g^{-1}(x)\)?

OpenStudy (anonymous):

4x

OpenStudy (jdoe0001):

ahemm, w0t?

OpenStudy (jdoe0001):

$$ g(x)=4-x \implies y=4-x\\ g^{-1} \implies x=4-y\\ \text{now solving by y}\\ x-4=-y \text{ multiply by -1 } \boldsymbol{-x+4=y}\\ g^{-1}(x)= ? $$